### v2

# Sensory Data Sharing in ISAC System: Fundamental Limits and Waveform Design

Abstract
The simultaneous acquisition and sharing of sensory data through a dual-functional signaling strategy, termed the communication-assisted sensing (CAS) system in this paper, has the potential to provide users with beyond-line-of-sight sensing capabilities. We mainly focus on three primary aspects, namely, the information-theoretic framework, the optimal distribution of channel input, and the optimal waveform design for Gaussian signals. First, we establish the information-theoretic framework and develop a modified source-channel separation theorem (MSST) tailored for CAS systems. The proposed MSST elucidates the relationship between achievable distortion, coding rate, and communication channel capacity in cases where the distortion metric is separable for sensing and communication (S&C) processes. Second, we present an optimal channel input design for dual-functional signaling, which aims to minimize CAS distortion under the constraints of the MSST and resource budget. We then conceive a two-step Blahut-Arimoto (BA)-based optimal search algorithm to numerically solve the functional optimization problem. Third, in light of the current signaling strategy, we further propose an optimal waveform design for Gaussian signaling in multi-input multi-output (MIMO) CAS systems. The associated covariance matrix optimization problem is addressed using a successive convex approximation (SCA)-based waveform design algorithm. Finally, we provide numerical simulation results to demonstrate the effectiveness of the proposed algorithms and to show the unique performance tradeoff between S&C processes.
Index Terms:
Integrated sensing and communication, communication assisted sensing, rate-distortion theory, source-channel separation theorem, Blahut-Arimoto algorithm.
I Introduction
Integrated sensing and communication (ISAC) systems are widely acknowledged for their potential to enhance sensing and communications (S&C) performance by sharing hardware and spectrum resources [2, 3]. Over the past few decades, substantial research efforts have focused on improving resource efficiency and mitigating mutual interference through waveform design [4], beamforming [5], and dedicated signal processing techniques [6]. To further explore the coordination gains between S&C, the authors of [7] proposed a sensing-assisted communication scheme, which leverages sensing capability of the transmitting ISAC signal, such as beam tracking and prediction, to establish user links thereby reducing the communication overhead. On the other hand, sensing-as-a-service is expected to become a key feature of upcoming 6G perceptive networks [8], which facilitates the reconsideration of how to share sensory measurement or target-related information (such as state variables) among devices or a fusion center within ISAC systems.
Figure 1: The application scenarios of the CAS system.
I-A Communication-Assisted Sensing Systems
The sharing of sensory data to extend coverage and enhance sensing accuracy has garnered extensive attention in traditional networks. For instance, the authors of [9] proposed a beyond-line-of-sight (BLoS) sensing scheme that shares measurements collected by vehicle-mounted sensors among users, enabling the detection of obstructed or distant targets. However, in ISAC systems, sensory data acquisition and sharing can be achieved simultaneously using a dual-functional signaling strategy, eliminating the need for additional sensors. This approach, referred to as the communication-assisted sensing (CAS) system in [10, 11], differs from traditional methods in that BLoS sensing is an intrinsic capability of the network.
Let us take the scenario in Fig. 1 as an example to illustrate the unique challenges in CAS sytems. The base station (BS) or roadside unit (RSU) with favorable visibility illuminates targets and captures observations during the sensing process while simultaneously transmitting the acquired sensory information to the end-users during the communication process. In a dual-functional signaling strategy, the ISAC waveform is used both to sense the current targets’ states and to convey the previously acquired state information to the end-users. This strategy inherently couples the S&C processes, giving rise to two unresolved issues in previous works.
1) The information theoretical framework of the CAS system has not been validated from a coding theory perspective, remaining an unclear operational meaning.
2) The ISAC waveform is assumed to be a Gaussian distribution in [11], which may not be necessarily optimal for the CAS scenario. Note the fact that Gaussian distribution is communication-optimal under Gaussian channels, whereas 2-ary pulse amplitude modulation is sensing-optimal [12].
This motivates us to delve into the fundamental limits and the optimal waveform design of the dual-functional signaling strategy for CAS systems.
Figure 2: The comparisons of the information-theoretic frameworks between the remote estimation, remote source coding, and the proposed CAS system.
I-B Related Works
Exploring the fundamental limits plays an important role in the ISAC systems [13]. The groundbreaking studies [12, 14] have considered a scenario where the transmitter (Tx) communicates with a user through a memoryless state-dependent channel while simultaneously estimating the state from generalized feedback. The capacity-distortion-cost tradeoff of this channel is characterized to illustrate the optimal achievable rate for reliable communication while maintaining a preset state estimation distortion. The work in [15] reveals the deterministic-random tradeoff between S&C within the dual-functional signaling strategy by characterizing the Cramér-Rao bound (CRB)-communication rate region. Unfortunately, the fundamental limits of CAS systems, whose setups significantly differ from the existing literature, remain widely unexplored.
The information-theoretic framework of the dual-functional signaling strategy for CAS systems is depicted in Fig 2, where $\mathsf{S}$, $\mathsf{Z}$, $\mathsf{X}$, $\mathsf{Y}$, and $\mathsf{\hat{S}}$ represent the source, the sensing channel observation, S&C channels input, the communication channel output, and the recovery at the receiver, respectively. Essentially, the CAS framework shares a similar methodology with conventional remote source coding (RSC), remote estimation (RE), and information bottleneck (IB) problems, by striving to balance the tradeoff between compression complexity (from $\mathsf{Z}$ to $\mathsf{\hat{S}}$) and the informative preservation (from $\mathsf{S}$ to $\mathsf{\hat{S}}$). However, we highlight the unique challenge introduced by adopting the ISAC signal $\mathsf{X}$ as both the S&C channels input.
$\bullet$ Remote source coding: In the RSC scheme, the noisy observation $\mathsf{Z}$ is encoded instead of the original source information $\mathsf{S}$ [16, 17, 18, 19, 20]. Rate-distortion theory is used to quantify the tradeoff between communication rate and achieved distortion, benefiting from the Markov chain $\mathsf{S}\leftrightarrow\mathsf{Z}\leftrightarrow\mathsf{X}\leftrightarrow\mathsf{Y}\leftrightarrow\mathsf{\hat{S}}$ [20]. However, in the CAS system, the observation $\mathsf{Z}$ depends on the ISAC channel input $\mathsf{X}$ due to the dual-functional signaling strategy. This dependency violates the Markov chain assumption, complicating the characterization of the rate-distortion relationship.
$\bullet$ Remote estimation: The RE involves a sensor measuring the state of a linear system and transmitting its observations to a remote estimator over a wireless fading channel [21, 22, 23, 24]. The sensory data-sharing scheme considered in [9] is a typical application in vehicular networks. This widely investigated problem differs from the CAS system in two ways. First, the RE requires additional sensors, whereas sensing capability is an intrinsic function in CAS system. Second, similar to the RSC, the observation $\mathsf{Z}$ and channel input $\mathsf{X}$ are decoupling.
$\bullet$ Information bottleneck: The IB problem is essentially equivalent to the RSC problem in which the distortion metric is specified by the logarithm loss fidelity criterion [25, 26]. For discussion convenience, we focus on the simplified process $\mathsf{S}\leftrightarrow\mathsf{Z}\leftrightarrow\mathsf{\hat{S}}$. The IB problem may be expressed by
$\mathop{\text{max}}\kern 2.0ptI(\mathsf{S};\mathsf{\hat{S}})\kern 5.0pt\text{subject to}\kern 5.0ptI(\mathsf{Z};\mathsf{\hat{S}})\leq\beta,$
where $I(\cdot,\cdot)$ denotes the mutual information (MI) and $\beta$ is the preset positive Lagrange multiplier that balances the tradeoff between compression and relevance. In contrast, when incorporating the ISAC signal $\mathsf{X}$ in CAS system, all variables in the IB problem become coupled. This leads to the following reformulation
[…]
where the tradeoff parameter $\beta$ is replaced by the communication channel capacity $I(\mathsf{X};\mathsf{Y})$ which also depends on $\mathsf{X}$. Evidently, the introduced ISAC signal significantly complicates the IB problem within the context of CAS system.
I-C Our contributions
Considering the aforementioned unique challenges, our aim is to establish the information-theoretic framework for the CAS systems, delving into the fundamental limits and specific waveform design. Compared to the conference version [1], which only discussed the corresponding information-theoretic framework, we further elaborate on the optimal distribution of channel input and propose a novel method for Gaussian waveform design in MIMO CAS systems. For clarity, the main contributions of this work are summarized as follows.
  • First, we establish the information-theoretic framework for CAS systems to illustrate the relationship between achievable distortion, coding rate, and communication channel capacity. We develop a modified source-channel separation theorem (MSST) specific to the cases of separable distortion metric for S&C processes. Compared to the existing works [11, 10], we provide a rigorous proof of the operational meaning for the MSST.
  • Second, we develop a optimal distribution design for the ISAC signal (S&C channel input), where the CAS distortion is minimized while adhering to the MSST and resource constraints. To cope with the functional optimization problem, we conceive a two-step Blahut-Arimoto (BA)-based optimal search algorithm in an effort to tackle the challenges of lacking explicit expressions for the rate-distortion bound and channel capacity.
  • Third, we propose an optimal waveform design scheme for Gaussian signaling in MIMO CAS systems. The associated covariance matrix optimization problem is solved using an successive convex approximation (SCA)-based waveform design algorithm.
  • Finally, we provide numerical simulation results to show the effectiveness of the proposed algorithms, while demonstrating the unique performance tradeoff between S&C in the considered CAS systems.
This paper is structured as follows. We commence with establishing the information-theoretic framework in Section II, including the definitions of distortions, rate-distortion function, and constrained channel capacity. In Section III, we prove the achievability and converse of the proposed MSST. Then, we formulate the optimization problem for ISAC channel input design and develop a two-step BA-based optimal search algorithm in Section IV. In Section V, we present the Gaussian ISAC waveform design method for the MIMO CAS systems. Finally, we provide the simulation results and conclude this paper in Section VI and VII, respectively.
The notations used in this paper are as follows. The uppercase normal letter $\mathsf{A}$, lowercase italic letter $a$, and fraktur letter $\mathcal{A}$ denote a random variable, its realization and a set, respectively. $P_{\mathsf{A}}(a)$ represent a probability distribution function and $Q_{\mathsf{A|B}}(a|b)$ specifies to channel transition probability. Uppercase and lowercase bold letters $\mathbf{A}$ and $\mathbf{a}$ denote the matrix and column vector. $(\cdot)^{T}$, $(\cdot)^{*}$, and $(\cdot)^{H}$ represent the transpose, conjugate, and complex conjugate transpose operations, respectively. $\mathbb{E}\left[\cdot\right]$ is the statistical expectation, and $\text{Tr}\left(\cdot\right)$ is the trace of a matrix.
II Information-theoretic framework of CAS
II-A Sensing and Communication Processes
As shown in Fig. 2, the random variables of the target’s state $\mathsf{S}$, the sensing observation $\mathsf{Z}$, the ISAC signal $\mathsf{X}$, and the communication channel output $\mathsf{Y}$ take values in the sets $\mathcal{S}$, $\mathcal{Z}$, $\mathcal{X}$, and $\mathcal{Y}$, respectively. Here, the state sequence $\{\mathsf{S}_{i}\}_{i\geq 1}$ is independently and identically distributed (i.i.d.) subject to a prior distribution $P_{\mathsf{S}}(s)$. In general, the CAS system may be described as the following S&C processes.
$\bullet$ Sensing Process: The sensing channel output, $\mathsf{Z}_{i}$, at a given time $i$ is generated based on the sensing channel law $Q_{\mathsf{Z|XS}}(\cdot|x_{i},s_{i})$ given the $i$th channel input $\mathsf{X}_{i}=x_{i}$ and the state realization $\mathsf{S}_{i}=s_{i}$. We assume that the observation $\mathsf{Z}_{i}$ is independent of past inputs, outputs and state signals.
$\bullet$ Communication Process: The ISAC signal $\mathsf{X}_{i}$ is obtained by encoding the observations $\mathsf{Z}^{i-1}$. The decoder receives the communication channel output $\mathsf{Y}_{i}$ in accordance with channel law $Q_{\mathsf{Y|X}}(y|x)$ and maps $\mathcal{Y}^{n}$ to $\hat{\mathcal{S}}^{n}$, where $\hat{\mathcal{S}}$ denotes the reconstruction alphabet.
By denoting $R$ as the bit rate, a $(2^{nR},n)$ coding scheme for the CAS systems consists of
1) A message set (quantized observations) $\mathcal{Z}^{n}=[1:2^{nR}]$;
2) A sequence of encoding functions $\phi_{i}$: $\mathcal{Z}^{i-1}\to\mathcal{X}$;
3) A decoder $\psi$: $\mathcal{Y}^{n}\to\hat{\mathcal{S}}^{n}$.
The performance of the CAS systems may be evaluated by the distortion between the ground truth and its reconstruction at the receiver. We term this distortion as the CAS distortion, which can be defined by
$\Delta^{(n)}:=\mathbb{E}\left[d(\mathsf{S}^{n},\hat{\mathsf{S}}^{n})\right]=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[d(\mathsf{S}_{i},\hat{\mathsf{S}}_{i})\right],$ (1)
where $d(\cdot,\cdot)$ is the distortion function bounded by $d_{\text{max}}$.
In practical systems, the S&C channel input $\mathsf{X}$ may be restricted by limited system resources. Let us define the cost-function $b(x):\mathcal{X}\to\mathbb{R}^{+}$ as the channel cost. Thus, a rate-distortion-cost tuple $(R,D,B)$ is said achievable if there exists a sequence of $(2^{nR},n)$ codes that simultaneously satisfy [14]
[…]
where $D$ and $B$ represent the CAS distortion and resource budget, respectively. We define the capacity-distortion-cost function $C_{\text{CAS}}(D,B)$ as the infimum of rate $R$ such that the tuple $(R,D,B)$ is achievable. As analyzed in Section I-B, compared to the existing works [12, 14], it is difficult to characterize such a capacity-distortion-cost curve due to the strong coupling between S&C process through the ISAC signal. In this paper, we characterize $C_{\text{CAS}}(D,B)$ within a estimate and compress (EC) strategy which achieves an optimal tradeoff between rate and distortion [27], and leave the general strategy for future research.
II-B The Estimate and Compress Strategy
In the EC strategy, the encoder transmits the estimate $\mathsf{\tilde{S}}$ extracted from the observation $\mathsf{Z}_{i}$, instead of transmitting $\mathsf{Z}$ itself. Therefore, the coding scheme may be recast by
1) A state parameter estimator $h$: $\mathcal{X}^{n}\times\mathcal{Z}^{n}\to\tilde{\mathcal{S}}^{n}$, where $\tilde{\mathcal{S}}$ denote the estimate alphabet;
2) A message set (quantized estimations) $\tilde{\mathcal{S}}^{n}=[1:2^{nR}]$;
3) A encoder function $\phi$: $\tilde{\mathcal{S}}^{n}\to\mathcal{X}^{n}$.
4) A decoder $\psi$: $\mathcal{Y}^{n}\to\hat{\mathcal{S}}^{n}$.
Although the estimator and encoder are defined in a block form for generality, the following Lemma 1 [14] show that the optimal estimator $h$ can be achieved by a symbol-by-symbol estimation.
Lemma 1 [14]: By recalling that $\mathsf{S}\leftrightarrow\mathsf{XZ}\leftrightarrow\mathsf{\tilde{S}}$ forms a Markov chain, the sensing distortion $\Delta_{s}^{(n)}$ is minimized by the deterministic estimator
$h^{\star}(x^{n},z^{n}):=(\tilde{s}^{\star}(x_{1},z_{1}),\tilde{s}^{\star}(x_{2},z_{2}),\cdots,\tilde{s}^{\star}(x_{n},z_{n})),$ (3)
where
$\tilde{s}^{\star}(x,z):=\arg\min_{s^{\prime}\in\mathcal{\tilde{S}}}\sum_{s\in\mathcal{S}}Q_{\mathsf{S|XZ}}(s|x,z)d(s,s^{\prime}),$ (4)
with posterior transition probability being
$Q_{\mathsf{S|XZ}}(s|x,z)=\frac{P_{\mathsf{S}}(s)Q_{\mathsf{Z|XS}}(z|x,s)}{\sum_{s^{\prime}\in\mathcal{S}}P_{\mathsf{S}}(s^{\prime})Q_{\mathsf{Z|XS}}(z|x,s^{\prime})},$
which is independent to the choice of encoder and decoder.
The detailed proof can be found in [14]. By applying the optimal estimator (4), the estimate $\mathsf{\tilde{S}}_{i}$ only depends on $\mathsf{X}_{i}$ and $\mathsf{Z}_{i}$ but independent of past inputs and outputs. Thus, the encoder of this scenario is reduced to a symbolwise encoding process. Moreover, by introducing the estimate $\mathsf{\tilde{S}}_{i}$, the CAS distortion $D$ can be divided into two parts, i.e., sensing distortion $D_{s}$ between $\mathsf{S}$ and $\mathsf{\tilde{S}}$ and communication distortion $D_{c}$ between $\mathsf{\tilde{S}}$ and $\mathsf{\hat{S}}$. Before presenting the MSST, we will give the definitions of $D_{s}$ and $D_{c}$.
In the sensing process, the estimation-cost function with respect to (w.r.t.) the ISAC channel input realization $\mathsf{X}=x$ can be defined by
$e(x)=\mathbb{E}\left[d(\mathsf{S},\tilde{s}^{\star}(\mathsf{X},\mathsf{Z}))|\mathsf{X}=x\right].$ (5)
Thus, a given sensing distortion $D_{s}$ satisfies
$\mathbb{E}\left[d(\mathsf{S},\mathsf{\tilde{S}})\right]=\mathbb{E}\left[e(\mathsf{X})\right]\leq D_{s}.$ (6)
In the communication process, the channel capacity depends on the ISAC channel input $\mathsf{X}$ which is constrained by a desired sensing distortion $D_{s}$ in the sensing process. This strong coupling property leads to a performance tradeoff between the sensing distortion and the achievable communication channel capacity. To proceed, we give the following definition.
Definition 1: The constrained channel capacity with sensing distortion $D_{s}$ and resource cost $B$ can be defined by
$C_{\text{IT}}(D_{s},B)=\mathop{\max}\limits_{P_{\mathsf{X}}(x)\in\mathcal{P}_{D_{s}}\cap\mathcal{P}_{B}}I(\mathsf{X};\mathsf{Y}),$ (7)
where $\mathcal{P}_{D_{s}}$ is the sensing feasible probability set whose element satisfies
$\mathcal{P}_{D_{s}}=\left\{P_{\mathsf{X}}(x)\big{|}\mathbb{E}\left[e(\mathsf{X})\right]\leq D_{s}\right\},$ (8)
and $\mathcal{P}_{B}$ is the resource feasible probability set described by
$\mathcal{P}_{B}=\left\{P_{\mathsf{X}}(x)\big{|}\mathbb{E}\left[b(\mathsf{X})\right]\leq B\right\}.$ (9)
The sequence $\{\mathsf{X}_{i}\}_{i\geq 1}$ is i.i.d. with distribution $P_{\mathsf{X}}(x)$ for achieving the capacity-distortion-cost region. By recalling the i.i.d. state sequence $\{\mathsf{S}_{i}\}_{i\geq 1}$, the observation $\{\mathsf{Z}_{i}\}_{i\geq 1}$ is also i.i.d. subject to the distribution
$P_{\mathsf{Z}}(z)=\sum_{s\in\mathcal{S}}\sum_{x\in\mathcal{X}}Q_{\mathsf{Z|XS}}(z|x,s)P_{\mathsf{S}}(s)P_{\mathsf{X}}(x).$ (10)
Furthermore, by adopting the symbol-by-symbol optimal estimator (4), the estimate $\{\tilde{\mathsf{S}}_{i}\}_{i\geq 1}$ is i.i.d. with the following distribution
$P_{\mathsf{\tilde{\mathsf{S}}}}(\tilde{\mathsf{s}})=\sum_{z\in\mathcal{Z}}\sum_{x\in\mathcal{X}}\sum_{s\in\mathcal{S}}Q_{\mathsf{Z|XS}}(z|x,s)P_{\mathsf{S}}(s)P_{\mathsf{X}}(x)\mathbb{I}\{\tilde{s}^{\star}(x,z)=\tilde{s}\},$ (11)
with $\mathbb{I}(\cdot)$ being the indicator function. Therefore, for the i.i.d. source $\mathsf{\tilde{S}}$, the relationship between $D_{c}$ and bit rate $R$ can be characterized by the rate-distortion theory.
Definition 2: The information-theoretic rate distortion function can be defined by
$R_{\text{IT}}(D_{c})=\mathop{\min}\limits_{P_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s}):\mathbb{E}\left[d(\mathsf{\tilde{S}},\mathsf{\hat{S}})\right]\leq D_{c}}I(\mathsf{\tilde{S}};\mathsf{\hat{S}}),$ (12)
where the minimization is over all conditional distributions $P_{\mathsf{\mathsf{\hat{S}}|\mathsf{\tilde{S}}}}(\hat{s}|\tilde{s})$ for which the joint distribution $P_{\mathsf{\tilde{S}\hat{S}}}(\tilde{s},\hat{s})=P_{\mathsf{\tilde{\mathsf{S}}}}(\tilde{\mathsf{s}})P_{\mathsf{\mathsf{\hat{S}}|\mathsf{\tilde{S}}}}(\hat{s}|\tilde{s})$ satisfies the expected distortion constraint. The rate-distortion function $R_{\text{IT}}(D_{c})$ demonstrates the minimal rates required to achieve a given communication distortion $D_{c}$.
III Modified Source-Channel Separation Theorem
In this section, we develop the modified source-channel separation theorem tailored for the dual-functional signaling CAS systems, aiming to elucidate the relationships between the CAS distortion, coding rate and communication channel capacity. The MSST retains a form analogous to the conventional source-channel separation theorem with distortion in lossy data transmission [28, Theorem 10.4.1].
Definition 3: The distortion metric $d(\cdot,\cdot)$ is separable for the S&C processes, if the following equality holds,
$\mathbb{E}\left[d(\mathsf{S},\mathsf{\hat{S}})\right]=\mathbb{E}\left[d(\mathsf{S},\mathsf{\tilde{S}})\right]+\mathbb{E}\left[d(\tilde{\mathsf{S}},\mathsf{\hat{S}})\right],$ (13)
implying that the CAS distortion $D$ is equal to $D_{s}+D_{c}$.
Note that the separability condition in Definition 3 appears to be strong. Fortunately, we demonstrate that the mean squared error (MSE) or quadratic distortion, a widely used distortion metric in parameter estimation, satisfies the separability condition. Specifically, we have
$D$ $=\mathbb{E}\left[\left\lVert\mathsf{S}-\hat{\mathsf{S}}\right\rVert_{2}^{2}\right]$ (14)
$\mathop{=}\limits^{(a)}\mathbb{E}\left[\left\|\mathsf{S}-\tilde{\mathsf{S}}\right\|_{2}^{2}\right]+\mathbb{E}\left[\left\|\tilde{\mathsf{S}}-\hat{\mathsf{S}}\right\|_{2}^{2}\right]\mathop{=}\limits^{\Delta}D_{s}+D_{c},$
where $(a)$ holds from the properties of the conditional expectation [20, Appendix A] with $\mathsf{\tilde{S}}=\mathbb{E}\left[\mathsf{S}|\mathsf{Z},\mathsf{X}\right]$ being the optimal estimator  [^1] We highlight that the (4) is exactly the minimum MSE (MMSE) estimator. An arbitrary non-optimal estimator except MMSE, which cannot apply the conditional expectation properties, may not satisfy the separability condition. . Now, we are ready to propose the MSST.
Theorem 1: The CAS distortion $D$ is achieved within a separable distortion metric, if and only if
$R_{\text{IT}}(D_{c})\leq C_{\text{IT}}(D_{s},B),$ (15)
where $C_{\text{IT}}(D_{s},B)$ and $R_{\text{IT}}(D_{c})$ are the constrained channel capacity (7) and the rate-distortion function (12), respectively.
The above MSST is proposed mathematically within the framework of information theory. Next, we elucidate the operational meaning of the MSST by proving that there must exist a practical coding scheme satisfying it.
III-A Converse
We start with a converse to show that any achievable coding scheme must satisfy (15). Consider a $(2^{nR},n)$ coding scheme defined by the encoding and decoding functions $\phi$ and $\psi$. Let $\mathsf{\tilde{S}}^{n}=h^{\star}(\mathsf{X}^{n},\mathsf{Z}^{n})$ be the estimate sequence as given in (3) and $\mathsf{\hat{S}}^{n}=\psi(\phi(\mathsf{\tilde{S}}^{n}))$ be the reconstruction sequence corresponding to $\mathsf{\tilde{S}}^{n}$.
Let us focus on the communication process $\mathsf{\tilde{S}}\leftrightarrow\mathsf{X}\leftrightarrow\mathsf{Y}\leftrightarrow\mathsf{\hat{S}}$. By recalling from the proof of the converse in lossy source coding, we have
$R$ $\geq\frac{1}{n}\sum_{i=1}^{n}I(\mathsf{\tilde{S}}_{i};\mathsf{\hat{S}}_{i})\mathop{\geq}\limits^{(a)}\frac{1}{n}\sum_{i=1}^{n}R_{\text{IT}}\left(\mathbb{E}\left[d(\mathsf{\tilde{S}}_{i},\mathsf{\hat{S}}_{i})\right]\right)$ (16)
$\mathop{\geq}\limits^{(b)}R_{\text{IT}}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[d(\mathsf{\tilde{S}}_{i},\mathsf{\hat{S}}_{i})\right]\Big{)}\mathop{\geq}\limits^{(c)}R_{\text{IT}}(D_{c}),$
where $(a)$ follows the Definition 2 that $R_{\text{IT}}$ is the minimum required MI, $(b)$ and $(c)$ are due to the convexity and non-increasing properties of the rate distortion function.
On the other hand, by recalling from the proof of the converse in channel coding, we have
$R$ $\leq\frac{1}{n}\sum_{i=1}^{n}I(\mathsf{X}_{i};\mathsf{Y}_{i})$ (17)
$\mathop{\leq}\limits^{(d)}\frac{1}{n}\sum_{i=1}^{n}C_{\text{IT}}\Big{(}\sum_{x\in\mathcal{X}}P_{\mathsf{X_{i}}}(x)c(x),\sum_{x\in\mathcal{X}}P_{\mathsf{X_{i}}}(x)b(x)\Big{)}$
$\mathop{\leq}\limits^{(e)}C_{\text{IT}}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\sum_{x\in\mathcal{X}}P_{\mathsf{X_{i}}}(x)c(x),\frac{1}{n}\sum_{i=1}^{n}\sum_{x\in\mathcal{X}}P_{\mathsf{X_{i}}}(x)b(x)\Big{)}$
$\mathop{\leq}\limits^{(f)}C_{\text{IT}}(D_{s},B),$
where $(d)$ follows the Definition 1 that the channel capacity is the maximum MI, $(e)$ and $(f)$ are due to the concavity and non-decreasing properties of the capacity constrained by estimation and resource costs [14]. By combing the inequalities (16), (17) and the data processing inequality $I(\mathsf{\tilde{S}};\mathsf{\hat{S}})\leq I(\mathsf{X};\mathsf{Y})$, we complete the proof of the converse.
III-B Achievability
Let $\mathsf{S}^{n}$ be drawn i.i.d. $\sim P_{\mathsf{S}}(s)$, we will show that there exists a coding scheme for a sufficiently large $n$ and rate $R$, the distortion $\Delta^{n}$ can be achieved by $D$ if (15) holds. The core idea follows the famous random coding argument and source channel separation theorem with distortion.
1) Codebook Generation: In source coding with rate distortion code, randomly generate a codebook $\mathcal{C}_{s}$ consisting of $2^{nR}$ sequences $\mathsf{\hat{S}}^{n}$ which is drawn i.i.d. $\sim P_{\mathsf{\hat{S}}}(\hat{s})$. The probability distribution is calculated by $P_{\mathsf{\hat{S}}}(\hat{s})=\sum_{\tilde{s}}P_{\mathsf{\tilde{S}}}(\tilde{s})Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$, where $P_{\mathsf{\tilde{S}}}(\tilde{s})$ is defined in (11) and $Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$ achieves the equality in (12). In channel coding, randomly generate a codebook $\mathcal{C}_{c}$ consisting of $2^{nR}$ sequences $\mathsf{X}^{n}$ which is drawn i.i.d. $\sim P_{\mathsf{X}}(x)$. The $P_{\mathsf{X}}(x)$ is chosen by satisfying the constrained capacity with estimation- and resource-cost in (7). Index the codeword $\mathsf{\tilde{S}}^{n}$ and $\mathsf{X}^{n}$ by $w\in\{1,2,\cdots,2^{nR}\}$.
2) Encoding: Encode the $\mathsf{\tilde{S}}^{n}$ by $w$ such that
$(\mathsf{\tilde{S}}^{n},\mathsf{\hat{S}}^{n}(w))\in\mathcal{T}^{(n)}_{d,\epsilon_{s}}(P_{\mathsf{\tilde{S}\hat{S}}}(\tilde{s},\hat{s})),$ (18)
where $\mathcal{T}^{(n)}_{d,\epsilon_{s}}(P_{\mathsf{\tilde{S}\hat{S}}}(\tilde{s},\hat{s}))$ represents the distortion typical set [28] with joint probability distribution $P_{\mathsf{\tilde{S}\hat{S}}}(\tilde{s},\hat{s})=P_{\mathsf{\tilde{S}}}(\tilde{s})Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$. To send the message $w$, the encoder transmits $x^{n}(w)$.
3) Decoding: The decoder observes the communication channel output $\mathsf{Y}^{n}=y^{n}$ and look for the index $\hat{w}$ such that
$(x^{n}(\hat{w}),y^{n})\in\mathcal{T}^{(n)}_{\epsilon_{c}}(P_{\mathsf{XY}}(x,y)),$ (19)
where $\mathcal{T}^{(n)}_{\epsilon_{c}}(P_{\mathsf{XY}}(x,y))$ represents the typical set with joint probability distribution $P_{\mathsf{XY}}(x,y)=P_{\mathsf{X}}(x)Q_{\mathsf{Y|X}}(y|x)$. If there exists such $\hat{w}$, it declares $\mathsf{\hat{S}}^{n}=\hat{s}^{n}(\hat{w})$. Otherwise, it declares an error.
4) Estimation: The encoder observes the channel output $\mathsf{Z}^{n}=z^{n}$, and computes the estimate sequence with the knowledge of channel input $x^{n}$ by using the estimator $\tilde{s}^{n}=h^{\star}(x^{n},z^{n})$ given in (3).
5) Distortion Analysis: We start by analyzing the expected communication distortion (averaged over the random codebooks, state and channel noise). In lossy source coding, for a fixed codebook $\mathcal{C}_{s}$ and choice of $\epsilon_{s}>0$, the sequence $\tilde{s}^{n}\in\tilde{\mathcal{S}}^{n}$ can be divided into two categories:
$\bullet$ $(\tilde{s}^{n},\hat{s}^{n}(w))\in\mathcal{T}^{(n)}_{d,\epsilon_{s}}$, we have $d(\tilde{s}^{n},\hat{s}^{n}(w))<D_{c}+\epsilon_{s}$;
$\bullet$ $(\tilde{s}^{n},\hat{s}^{n}(w))\notin\mathcal{T}^{(n)}_{d,\epsilon_{s}}$, we denote $P_{e_{s}}$ as the total probability of these sequences. Thus, these sequence contribute at most $P_{e_{s}}d_{\max}$ to the expected distortion since the distortion for any individual sequence is bounded by $d_{\max}$.
According to the achievability of lossy source coding [28, Theorem 10.2.1], we have $P_{e_{s}}$ tends to zero for sufficiently large $n$ whenever $R\geq R_{\text{IT}}(D_{c})$.
In channel coding, the decoder declares an error when the following events occur:
$\bullet$ $(x^{n}(w),y^{n})\notin\mathcal{T}^{(n)}_{\epsilon_{c}}$;
$\bullet$ $(x^{n}(w^{\prime}),y^{n})\in\mathcal{T}^{(n)}_{\epsilon_{c}}$, for some $w^{\prime}\neq w$, we denote $P_{e_{c}}$ as the probability of the error occurred in decoder. The error decoding contribute at most $P_{e_{c}}d_{\max}$ to the expected distortion. Similarly, we have $P_{e_{c}}\to 0$ for $n\to\infty$ whenever $R\leq C_{\text{IT}}(D_{s},B)$ according to channel coding theorem [28].
On the other hand, the expected estimation distortion can be upper bounded by
$\Delta_{s}^{(n)}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[d(\mathsf{S}_{i},\mathsf{\tilde{S}}_{i})|\mathsf{\hat{W}}\neq w\right]\text{Pr}(\mathsf{\hat{W}}\neq w)$ (20)
$\kern 30.0pt+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[d(\mathsf{S}_{i},\mathsf{\tilde{S}}_{i})|\mathsf{\hat{W}}=w\right]\text{Pr}(\mathsf{\hat{W}}=w)$
$\leq P_{e_{c}}d_{\text{max}}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[d(\mathsf{S}_{i},\mathsf{\tilde{S}}_{i})|\mathsf{\hat{W}}=w\right](1-P_{e_{c}}).$
Note that $(s^{n},x^{n}(w),\tilde{s}^{n})\in\mathcal{T}^{(n)}_{\epsilon_{e}}(P_{\mathsf{SX\tilde{S}}}(s,x,\tilde{s}))$ where $P_{\mathsf{SX\tilde{S}}}(s,x,\tilde{s})$ denotes the joint marginal distribution of $P_{\mathsf{SXZ\tilde{S}}}(s,x,z,\tilde{s})=P_{\mathsf{S}}(s)P_{\mathsf{X}}(x)Q_{\mathsf{Z|SX}}(z|s,x)\mathbb{I}\{\tilde{s}=\tilde{s}^{\star}(x,z)\}$, we have
[…]
according to the typical average lemma [14]. In summary, the CAS distortion can be attained by
$\Delta^{(n)}$ $\mathop{=}\limits^{(a)}\Delta_{s}^{(n)}+\Delta_{c}^{(n)}$ (22)
$\mathop{\leq}\limits^{(b)}D_{c}+(P_{e_{s}}+2P_{e_{c}})d_{\max}+(1+\epsilon_{e})(1-P_{e_{c}})D_{s},$
where $(a)$ follows the condition of the separable distortion metric in MSST and we omit the terms containing the product of $P_{e_{s}}$ and $P_{e_{c}}$ in step $(b)$. Consequently, taking $n\to\infty$ and $P_{e_{s}},P_{e_{c}},\epsilon_{c},\epsilon_{e},\epsilon_{s}\to 0$, we can conclude that the expected CAS distortion (averaged over the random codebooks, state and channel noise) tends to be $D\to D_{c}+D_{s}$ whenever $R_{\text{IT}}(D_{c})\leq R\leq C_{\text{IT}}(D_{s},B)$.
[…]
Remark: The sensing process determines the accuracy of the state information acquisition at the encoder, whereas the communication process governs the quantity of information transmitted to the decoder. Overemphasis on either process can result in substantial performance degradation. For instance, achieving optimal sensing performance alone may not ensure the accurate reconstruction at the decoder due to limitations in the communication channel capacity. Conversely, excessive communication capacity may be wasted if the sensory data lacks sufficient accuracy. This underscores the need to investigate optimal channel input as well as the waveform design that minimize the CAS distortion.
IV Channel Input Distribution Optimization for the CAS Systems
In this section, we explore the optimal input distribution for the CAS system under scalar channels, aiming to minimize the CAS distortion under the MMST constraint. Hereinafter, we adopt the MSE as the distortion metric to guarantee the MSST. Meanwhile, we drop the subscript “IT” in the MSST.
IV-A Problem Formulation
The ISAC channel input design can be formulated by the following optimization problem  [^2] Here, we use the symbol $I(D_{s},B)$ instead of $C(D_{s},B)$ to emphasize that the optimal distribution does not necessarily achieve the channel capacity.
$\mathcal{P}_{0}\left\{\kern 2.0pt\begin{aligned} \mathop{\text{min}}\limits_{P_{\mathsf{X}}(x)}\kern 5.0pt&D=D_{c}+D_{s}\\ \text{subject to}\kern 5.0pt&R(D_{c})\leq I(D_{s},B).\end{aligned}\right.$ (23)
By substituting the associated expressions into (23), problem $\mathcal{P}_{0}$ can be reformulated by
$\mathcal{P}_{1}\left\{\kern 2.0pt\begin{aligned} \mathop{\text{min}}\limits_{P_{\mathsf{X}}(x)}\kern 5.0pt&\mathbb{E}\left[d(\mathsf{S},\tilde{\mathsf{S}})\right]+\mathbb{E}\left[d(\mathsf{\tilde{S}},\hat{\mathsf{S}})\right]\\ \text{subject to}\kern 5.0pt&I(\mathsf{\tilde{S};\hat{S}})\leq I(\mathsf{X};\mathsf{Y}),\\ \kern 5.0pt&\mathbb{E}\left[b(\mathsf{X})\right]\leq B.\end{aligned}\right.$ (24)
It is challenging to directly solve the functional optimization problem $\mathcal{P}_{1}$ due to the difficulty in obtaining explicit expressions of distortions and MI for arbitrary distributions. Typically, the BA algorithm can be employed to solve problem $\mathcal{P}_{1}$ numerically. Inspired by this idea, we expand the expectation operation and the MI w.r.t. the variable $P_{\mathsf{X}}(x)$ as follows.
$\bullet$ Sensing distortion:
$\mathbb{E}\left[d(\mathsf{S},\mathsf{\tilde{S}})\right]$ $=\mathbb{E}\left[\mathbb{E}\left[d(\mathsf{S},\tilde{s}(\mathsf{X},\mathsf{Z}))|\mathsf{X},\mathsf{Z}\right]\right]$ (25)
$=\sum_{x,z}P_{\mathsf{X}\mathsf{Z}}(x,z)\sum_{s}Q_{\mathsf{S}|\mathsf{X}\mathsf{Z}}(s|x,z)d(s,\tilde{s}(x,z))$
$\triangleq\sum_{x}P_{\mathsf{X}}(x)e(x),$
where the sensing cost $e(x)$ defined in (5) can be expressed by
$e(x)=\sum_{z}Q_{\mathsf{Z}|\mathsf{X}}(z|x)\sum_{s}Q_{\mathsf{S}|\mathsf{X}\mathsf{Z}}(s|x,z)d(s,\tilde{s}(x,z)).$ (26)
$\bullet$ Channel MI:
$I(\mathsf{X;Y})=\sum_{x,y}P_{\mathsf{X}}(x)Q_{\mathsf{Y|X}}(y|x)\log\frac{Q_{\mathsf{X|Y}}(x|y)}{P_{\mathsf{X}}(x)}.$ (27)
$\bullet$ Resource budget:
$\mathbb{E}\left[b(\mathsf{X})\right]=\sum_{x}P_{\mathsf{X}}(x)b(x).$ (28)
$\bullet$ Communication distortion:
$\mathbb{E}[d(\mathsf{\tilde{S},\hat{S}})]=\sum_{\tilde{s}}P_{\mathsf{\tilde{S}}}(\tilde{s})\sum_{\hat{s}}Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})d(\tilde{s},\hat{s}).$ (29)
$\bullet$ Rate distortion function:
$I(\mathsf{\tilde{S};\hat{S}})=\sum_{\tilde{s},\hat{s}}P_{\mathsf{\tilde{S}}}(\tilde{s})Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})\log\frac{Q_{\mathsf{\tilde{S}|\hat{S}}}(\tilde{s}|\hat{s})}{P_{\mathsf{\tilde{S}}}(\tilde{s})}.$ (30)
Here, we emphasize the unique challenge in solving problem $\mathcal{P}_{1}$. In the conventional BA algorithm, the primary approach involves constructing the Lagrangian function w.r.t. the variable and obtaining the optimal solution based on the first-order necessary condition. However, the probability distribution $P_{\mathsf{\tilde{S}}}(\tilde{s})$ in (29) and (30) is indeed a function of the variable $P_{\mathsf{X}}(x)$ as describe in (11). Consequently, deriving a closed-form expression of $P_{\mathsf{X}}(x)$ from first-order necessary condition is challenging, which imposes a significant obstacle to solving problem $\mathcal{P}_{1}$ by using the conventional BA algorithm. To address this issue, we propose a two-step BA-based optimal search method to seek for a sub-optimal solution in the following subsection.
IV-B Two-step BA-based Optimal Search Algorithm
We divide the original problem $\mathcal{P}_{1}$ into two sub-problems.
1) Sub-problem 1: We determine the constrained communication channel capacity $I(\mathsf{X},\mathsf{Y})$ for a given sensing distortion $D_{s}$ and identify the optimal distribution $P_{\mathsf{X}}(x)$.
2) Sub-problem 2: We calculate the minimum communication distortion $D_{c}$ for the given channel input distribution $P_{\mathsf{X}}(x)$ and determine the communication distortion $D_{c}$, thereby obtaining the CAS distortion $D$.
By varying the preset sensing distortions, one may generate a set $\mathcal{D}$ collecting the CAS distortions calculated through the above two steps. Thus, the minimum CAS distortion and its corresponding distribution $P_{\mathsf{X}}(x)$ can be identified by the minimum value search. The detailed procedure is as follows.
$\bullet$ Constrained communication channel Capacity for a given sensing distortion.
In this sub-problem, $I(\mathsf{X};\mathsf{Y})$ is maximized under the constraints of sensing distortion $D_{s}$ and resource budget $B$, which can be expressed by
$\mathop{\text{max}}\limits_{P_{\mathsf{X}}(x)}\kern 10.0pt$ $I(\mathsf{X};\mathsf{Y})$ (31)
subject to $\kern 5.0pt\sum_{x}P_{\mathsf{X}}(x)e(x)\leq D_{s},\kern 5.0pt\sum_{x}P_{\mathsf{X}}(x)b(x)\leq B.$
Inspired by [12], we reformulate (31) by incorporating the sensing distortion cost as a penalty term in the objective function, i.e.,
$\mathop{\text{max}}\limits_{P_{\mathsf{X}}(x)}\kern 5.0pt$ $I(P_{\mathsf{X}}(x),Q_{\mathsf{Y|X}}(y|x))-\mu\sum_{x}P_{\mathsf{X}}(x)e(x)$ (32)
subject to $\kern 5.0pt\sum_{x}P_{\mathsf{X}}(x)b(x)\leq B,$
where $\mu$ is the penalty factor to balance the weight of S&C performance. For a given $\mu$, the Lagrangian function of (32) can be written by
$\mathcal{L}(P_{\mathsf{X}}(x),\lambda)$ $=I(P_{\mathsf{X}}(x),Q_{\mathsf{Y|X}}(y|x))$ (33)
$-\mu\sum_{x}P_{\mathsf{X}}(x)e(x)+\lambda\sum_{x}P_{\mathsf{X}}(x)b(x).$
By setting the derivative $\partial\mathcal{L}(P_{\mathsf{X}}(x),\lambda)/\partial P_{\mathsf{X}}(x)$ to zero and taking the fact that $\sum_{x}P_{\mathsf{X}}(x)=1$ into account, we have
$P_{\mathsf{X}}(x)=\frac{e^{h(x)}}{\sum_{x^{\prime}}e^{h(x^{\prime})}},$ (34)
with
$h(x)=\sum_{y}Q_{\mathsf{Y|X}}(y|x)\log Q_{\mathsf{X|Y}}(x|y)-\mu e(x)-\lambda b(x).$ (35)
In formula (34), the Lagrangian multiplier $\lambda$ and the posterior distribution $Q_{\mathsf{X|Y}}(x|y)$ are unknown. The former may be determined by a bisection search method such that $\sum_{x}P_{\mathsf{X}}(x)b(x)\to B$. For the latter, we have the optimal posterior distribution which maximizes the MI with given $P_{\mathsf{X}}(x)$ and $Q_{\mathsf{Y|X}}(y|x)$, expressed by [29, Lemma 9.1]
$Q_{\mathsf{X|Y}}(x|y)=\frac{P_{\mathsf{X}}(x)Q_{\mathsf{Y|X}}(y|x)}{\sum_{x^{\prime}}P_{\mathsf{X}}(x^{\prime})Q_{\mathsf{Y|X}}(y|x^{\prime})}.$ (36)
After setting an initial distribution $P_{\mathsf{X_{0}}}(x)$, problem (32) can be solved by updating (36) and (34) iteratively. The detailed procedure is summarized in Algorithm 1, whose convergence can be guaranteed [14, 30].
Input: $P_{\mathsf{S}}(s)$, $Q_{\mathsf{Z|SX}}(z|s,x)$ and $Q_{\mathsf{Y|X}}(y|x)$, $B$, $\mu$, $K$.
1 Initialize: Set $P_{\mathsf{X_{0}}}(x)$ as uniform distribution;
2 for $k\leq K$ do
3       Step 1: Compute sensing cost $e(x)$ in (4) ;
4       Step 2: Update posterior distribution in (36) ;
5       Step 3: Update $P_{\mathsf{X}}(x)$ in (34) ;
6       Step 4: Find $\lambda^{\star}$ by bisection searching ;
7
8 end for
Output: Optimal $P_{\mathsf{X}}(x)$, $D_{s}$ and $I(\mathsf{X};\mathsf{Y})$.
Algorithm 1 Modified Iteration BA Algorithm
$\bullet$ Source distribution $P_{\mathsf{\tilde{S}}}(\tilde{s})$.
The tradeoff between sensing distortion and constrained communication capacity can be adjusted by varying the penalty factor $\mu$. Let us denote the optimal channel input distribution, sensing distortion, and channel MI obtained by Algorithm 1 as $P_{\mathsf{X}}^{(\mu)}(x)$, $D_{s}^{(\mu)}$, and $I^{(\mu)}(\mathsf{X};\mathsf{Y})$, respectively, to highlight their dependence on the factor $\mu$.
Subsequently, given a specific channel input distribution $P_{\mathsf{X}}^{(\mu)}(x)$, our objective is to compute the estimate distribution $P_{\mathsf{\tilde{S}}}^{(\mu)}(\tilde{s})$, which is essential for determining the rate-distortion function. However, deriving the explicit expression of the estimate distribution for an arbitrary state distribution $P_{\mathsf{S}}(s)$ is a challenging task. Instead, we can compute the estimate $\tilde{\mathsf{S}}$ and its distribution $P_{\mathsf{\tilde{S}}}^{(\mu)}(\tilde{s})$ using formulas (4) and (11), respectively, by performing a sufficiently large number of random trials. For analytical convenience, we consider a Gaussian state distribution and a linear sensing model, which allows for explicit expressions to be derived.
Assume that the target’s state follows Gaussian distribution with $\mathcal{CN}(0,\nu_{s}^{2})$. Let us consider the following linear sensing model
$\mathsf{Z}=\mathsf{XS}+\mathsf{N},$ (37)
where $\mathsf{N}\sim\mathcal{CN}(0,1)$ represents Gaussian channel noise. In such a Gaussian linear model, the estimator (4) is indeed an MMSE estimator for each channel realization $x$. Therefore, the estimate $\mathsf{\tilde{S}}$ and the corresponding conditional probability distribution of $P_{\mathsf{\tilde{S}|X}}(\tilde{s}|x)$ can be expressed by
$\mathsf{\tilde{S}}=\frac{x^{2}\nu_{s}^{4}}{(1+x^{2}\nu_{s}^{2})^{2}}\mathsf{Z},\kern 5.0ptP_{\mathsf{\tilde{S}|X}}(\tilde{s}|x)\sim\mathcal{CN}(0,\frac{x^{2}\nu_{s}^{4}}{1+x^{2}\nu_{s}^{2}}).$ (38)
Moreover, sensing cost is the MSE with the given realization $x$, which can be written by
$e(x)=\frac{\nu_{s}^{2}}{(1+x^{2}\nu_{s}^{2})}.$ (39)
Consequently, the source distribution can be given by
$P_{\mathsf{\tilde{S}}}^{(\mu)}(\tilde{s})=P_{\mathsf{X}}^{(\mu)}(x)P_{\mathsf{\tilde{S}|X}}(\tilde{s}|x).$ (40)
$\bullet$ Communication distortion.
In this sub-problem, we aim to evaluate the minimum communication distortion $D_{c}^{(\mu)}$ achieved under the source distribution $P_{\mathsf{\tilde{S}}}^{(\mu)}(\tilde{s})$ and the channel capacity $I^{(\mu)}(\mathsf{X};\mathsf{Y})$. This scenario corresponds to a typical rate-distortion problem in lossy data transmission. The distortion-rate curve can be obtained by the optimization problem [29, Chapter 9]
$\mathop{\text{min}}\limits_{Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})}\kern 5.0ptI(\mathsf{\tilde{S};\hat{S}})-\lambda_{s}D_{c}$ (41)
where $\lambda_{s}$ represents the slope of the distortion-rate curve. The conventional BA algorithm can be leveraged to solve problem (41). By following a similar procedure in sub-problem 1, the optimal solution can be calculated by
$Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})=\frac{P_{\mathsf{\hat{S}}}(\hat{s})e^{\lambda_{s}d(\tilde{s},\hat{s})}}{\sum_{\hat{s}}P_{\mathsf{\hat{S}}}(\hat{s}^{\prime})e^{\lambda_{s}d(\tilde{s},\hat{s}^{\prime})}}.$ (42)
The unknown nuisance term $P_{\mathsf{\hat{S}}}(\hat{s})$ can be updated by
$P_{\mathsf{\hat{S}}}(\hat{s})=\sum_{\tilde{s}}P_{\mathsf{\tilde{S}}}(\tilde{s})Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s}).$ (43)
By setting an initial distribution $Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$, the optimal solution of (41) can be obtained by updating (43) and (42) iteratively. The detailed derivations can be found in [29].
For a given slop $\lambda_{s}$, the tangent point $\big{(}D^{(\lambda_{s})}_{c},I^{(\lambda_{s})}(\mathsf{\tilde{S};\hat{S}})\big{)}$ of the distortion-rate curve can be obtained by solving problem (41). We may collect a series of distortion-rate data points to fit a distortion-rate function by varying the slop $\lambda_{s}$. Thus, the communication distortion constrained by a given channel capacity is attained by
$D_{c}^{(u)}=D_{\text{IT}}(I^{(u)}(\mathsf{X};\mathsf{Y})),$ (44)
where $D_{\text{IT}}(\cdot)$ represents the distortion-rate function. We summarize the communication distortion computation algorithm into Algorithm 2.
$\bullet$ CAS distortion.
In the outer loop, we may collect a set $\mathcal{D}$ of the CAS distortion $D^{(\mu)}=D_{s}^{(u)}+D_{c}^{(u)}$ and the associated channel input distribution $P_{\mathsf{X}}^{(\mu)}(x)$ by repeating the above steps with various penalty factor $\mu$. Therefore, the optimal channel distribution that minimizes the CAS distortion can be found by
[…]
The proposed two-step BA-based optimal search algorithm is summarized in Algorithm 3.
Input: $P_{\mathsf{\tilde{S}}}(\tilde{s})$, $I^{(u)}(\mathsf{X};\mathsf{Y})$, $K$, a set of the slops $\mathcal{A}$.
1 for $\lambda_{s}\in\mathcal{A}$ do
2       Initialize: Set $Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$ as uniform distribution ;
3       for $k\leq K$ do
4             Step 1: Update output distritbution in (43) ;
5             Step 2: Update $Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$ in (42) ;
6
7       end for
8      Step 3: Collect data $\mathcal{I}\leftarrow\big{(}D_{c}^{(\lambda_{s})},I^{(\lambda_{s})}(\mathsf{\tilde{S};\hat{S}})\big{)}$ ;
9
10 end for
11Step 4: Characterize the $D(R)$ function base on $\mathcal{I}$ ;
Output: Communication distortion in (44).
Algorithm 2 Communication Distortion Algorithm
Input: The set of penalty factor $\mathcal{U}\subset[0,\mu_{\text{max}}]$.
1 for $u\in\mathcal{U}$ do
2       Step 1: Obtain $P^{(\mu)}_{\mathsf{X}}(x)$, $D^{(\mu)_{s}}$, and $I^{(u)}(\mathsf{X};\mathsf{Y})$ through Algorithm 1 ;
3       Step 2: Calculate estimate distribution $P^{(\mu)}_{\mathsf{\tilde{S}}}(\tilde{s})$ ;
4       Step 3: Obtain $D_{c}^{(\mu)}$ through Algorithm 2 ;
5       Step 4: $\mathcal{D}\leftarrow D^{(\mu)}=D_{s}^{(\mu)}+D_{c}^{(\mu)}$ ;
6
7 end for
Output: Optimal solution in (45).
Algorithm 3 Two-step BA-based Optimal Search
V Waveform Design for Gaussian Channel Input
The previous section illustrates that using Gaussian channel input may not be optimal for the CAS systems. Despite this limitation, the assumption of Gaussian signals is prevalent in communications research literature. In this section, we introduce an ISAC waveform design scheme specifically developed for multi-input multi-output (MIMO) systems, while adhering to the framework of Gaussian signaling.
V-A System Model
The widely employed signal models of the S&C processes for the MIMO systems are expressed by
$\mathbf{Z}=\mathbf{H}_{s}\mathbf{X}+\mathbf{N}_{s},\kern 5.0pt\mathbf{Y}=\mathbf{H}_{c}\mathbf{X}+\mathbf{N}_{c},$ (46)
where $\mathbf{X}\in\mathbb{C}^{N_{t}\times T}$, $\mathbf{Z}\in\mathbb{C}^{M_{s}\times T}$, and $\mathbf{Y}\in\mathbb{C}^{M_{c}\times T}$ represent the transmitting ISAC signal, the received signals at the BS (encoder) and the end-user (decoder), respectively; $\mathbf{H}_{s}$ and $\mathbf{H}_{c}$ denote the S&C channel state information (CSI) matrices; $\mathbf{N}_{s}$ and $\mathbf{N}_{c}$ are the channel noises whose entries follow the complex Gaussian distribution with $\mathcal{CN}(0,\sigma_{s}^{2})$ and $\mathcal{CN}(0,\sigma_{c}^{2})$; Finally, $N_{t}$, $M_{s}$, $M_{c}$, and $T$ denote the numbers of Tx antennas, S&C receiver (Rx) antennas, and the transmitting symbols, respectively.
We focus on the sensing task of target response matrix (TRM) estimation, where the to-be-estimate parameters can be written by $\mathbf{s}=\text{vec}(\mathbf{H}_{s})$. After vectorization operations, the sensing signal observed at the BS can be recast by
$\mathbf{z}=\left(\mathbf{I}_{M_{s}}\otimes\mathbf{X}^{H}\right)\mathbf{s}+\mathbf{n}_{s},$ (47)
with $\mathbf{z}=\text{vec}(\mathbf{Z}^{H})$, and $\mathbf{n}_{s}=\text{vec}(\mathbf{N}_{s}^{H})$. The BS estimates TRM $\tilde{\mathbf{s}}$ from the observations $\mathbf{z}$, then transmits it to the user through communication channel $\mathbf{H}_{c}$. Subsequently, the user can reconstruct $\hat{\mathbf{s}}$ from the communication received data $\mathbf{Y}$. Here, we adopt the quadratic distortion metric (MSE), which meets the separability condition as shown in (14). Before proceeding the ISAC waveform design, we make the following assumptions.
Assumption 1: The ISAC waveform $\mathbf{X}$ follows complex Gaussian distribution with $\mathcal{CN}(0,\mathbf{R}_{x})$, where $\mathbf{R}_{x}\in\mathbb{C}^{N_{t}\times N_{t}}$ is the covariance matrix.
Assumption 2: The TRM vector $\mathbf{s}$ follows complex Gaussian distribution $\mathcal{CN}(0,\mathbf{I}_{M_{s}}\otimes\bm{\Sigma}_{s})$, where $\bm{\Sigma}_{s}\in\mathbb{C}^{N_{t}\times N_{t}}$ denotes the covariance matrix of each column of $\mathbf{H}_{s}$. [^3] This assumption corresponds to the scenario that the Rx antennas for sensing are sufficiently separated so that the correlations among the rows of $\mathbf{H}_{s}$ can be ignored [31]. Here, we specify this Kronecker structure in TRM covariance matrix to simplify the expression of sensing distortion [32]. To avoid the deviation of our core contribution in this paper, we will leave the general TRM covariance matrix cases for future research.
The above assumptions enable us to derive the explicit expressions of the S&C distortions, the estimate distribution $\tilde{\mathbf{s}}$, and the communication channel capacity w.r.t. the ISAC waveform matrix $\mathbf{X}$.
1) Sensing distortion: For the Gaussian linear model (47), the optimal estimator (4), i.e., the well-known MMSE estimator, is leveraged to obtain the estimate $\tilde{\mathbf{s}}$ by
$\tilde{\mathbf{s}}=\Big{(}\mathbf{I}_{M_{s}}\otimes\Big{(}\bm{\Sigma}_{s}\mathbf{X}\left(\mathbf{X}^{H}\bm{\Sigma}_{s}\mathbf{X}+\sigma^{2}_{s}\mathbf{I}_{T}\right)^{-1}\Big{)}\Big{)}\mathbf{z},$ (48)
Furthermore, the sensing distortion is attained by [33]
$D_{s}(\mathbf{R}_{x})=\mathbb{E}\Big{[}\|\mathbf{s}-\tilde{\mathbf{s}}\|^{2}\Big{]}=M_{s}\text{Tr}\left(\Big{(}\frac{1}{\sigma^{2}_{s}}\mathbf{R}_{x}+\bm{\Sigma}_{s}^{-1}\Big{)}^{-1}\right).$ (49)
2) Communication distortion: The estimate $\tilde{\mathbf{s}}$ in (48) follows the complex Gaussian distribution with $\mathcal{CN}(\mathbf{0},\mathbf{I}_{M_{s}}\otimes\mathbf{R}_{\tilde{s}})$, with the covariance matrix of [33]
$\mathbf{R}_{\tilde{\mathbf{s}}}=\bm{\Sigma}_{s}-\Big{(}\frac{1}{\sigma^{2}_{s}}\mathbf{R}_{x}+\bm{\Sigma}_{s}^{-1}\Big{)}^{-1}.$ (50)
Let us temporarily omit the Kronecker product. For a Gaussian source $\mathcal{CN}(\mathbf{0},\mathbf{R}_{\tilde{s}})$, the rate-distortion function can be characterized by the following optimization problem [34]
$R(D_{c})=\mathop{\min}\limits_{\mathbf{D}\in\mathbb{S}_{N_{t}}^{+}}$ $\log\left(\frac{\det(\mathbf{R}_{\tilde{s}})}{\det(\mathbf{D})}\right)$ (51)
subject to $\kern 5.0pt\mathbf{D}\preceq\mathbf{R}_{\tilde{s}},\kern 2.0pt\text{Tr}\left(\mathbf{D}\right)\leq D_{c},$
where $\mathbb{S}_{N_{t}}^{+}$ denotes the set of all $N_{t}\times N_{t}$ positive definite matrices. More details on (51) can be found in the proof of [34, Theorem 2].
3) Channel Capacity: The MIMO Gaussian channel capacity can be expressed by
$C(\mathbf{R}_{x})=I(\mathbf{Y};\mathbf{X}|\mathbf{H}_{c})=\log\det\left(\frac{1}{\sigma^{2}_{c}}\mathbf{H}_{c}\mathbf{R}_{x}\mathbf{H}_{c}^{H}+\mathbf{I}_{M_{c}}\right).$ (52)
Input: An initial covariance matrix $\mathbf{R}_{0}$, $K$.
1 for $k\leq K$ do
2       Step 1: Obtain optimal $\mathbf{R}^{\star}_{x}$ by solving $\mathcal{P}_{3}$ ;
3       Step 2: $\mathbf{R}_{0}=\mathbf{R}^{\star}_{x}$ ;
4
5 end for
Output: Optimal solution $\mathbf{R}^{\star}_{x}$.
Algorithm 4 SCA-based Waveform Design Algorithm
V-B ISAC Waveform Design
Observe that the expressions (49) to (52) are all related to the covariance matrix $\mathbf{R}_{x}$ rather than the specific waveform matrix $\mathbf{X}$. Consequently, our ISAC waveform design problem can be transformed into determining the optimal covariance matrix that minimizes the CAS distortion. Regarding the system resource constraint, we only consider the transmit power at the BS. By substituting the expressions into the original problem $\mathcal{P}_{1}$ and introducing an auxiliary variable of positive definite matrix $\mathbf{D}$, the ISAC waveform design problem can be reformulated by
$\mathcal{P}_{2}\left\{\begin{aligned} \mathop{\min}\limits_{\mathbf{R}_{x},\mathbf{D}\in\mathbb{S}_{N_{t}}^{+}}&D_{s}(\mathbf{R}_{x})+\text{Tr}\left(\mathbf{D}\right)\\ \text{subject to}&\kern 5.0ptC(\mathbf{R}_{x})-M_{s}\log\left(\frac{\det(\mathbf{R}_{\tilde{s}})}{\det(\mathbf{D})}\right)\geq 0,\\ &\kern 5.0pt\mathbf{D}\preceq\mathbf{R}_{\tilde{s}},\kern 2.0pt\mathbf{R}_{x}\succeq 0,\kern 2.0pt\text{Tr}\left(\mathbf{R}_{x}\right)\leq P_{T},\end{aligned}\right.$ (53)
where $P_{T}$ is the power budget. $\mathcal{P}_{2}$ is non-convex due to the nonlinear intermediate variables $\log\det(\mathbf{R}_{\tilde{s}})$ and $\mathbf{R}_{\tilde{s}}$. To relax $\mathcal{P}_{2}$ into a convex problem, we employ an SCA technique based on Taylor series expansion. To be specified, for any given point $\mathbf{R}_{0}$, the matrix $\mathbf{R}_{\tilde{s}}$ can be approximated by
$\mathbf{R}_{\tilde{s}}\simeq\bm{\Sigma}_{s}-\mathbf{P}+\frac{1}{\sigma_{s}}\mathbf{P}(\mathbf{R}_{x}-\mathbf{R}_{0})\mathbf{P}\triangleq\widetilde{\mathbf{R}}_{\tilde{s}},$ (54)
with the constant matrix defined by
$\mathbf{P}=\Big{(}\frac{1}{\sigma^{2}_{s}}\mathbf{R}_{0}+\bm{\Sigma}_{s}^{-1}\Big{)}^{-1}.$
Furthermore, we have
$\log\det(\mathbf{R}_{\tilde{s}})\simeq\log\det\left(\bm{\Sigma}_{s}-\mathbf{P}\right)$ (55)
$\kern 10.0pt+\text{Tr}\left(\frac{1}{\sigma_{s}^{2}}(\bm{\Sigma}_{s}-\mathbf{P})^{-1}\mathbf{P}(\mathbf{R}_{x}-\mathbf{R}_{0})\mathbf{P}\right)\triangleq f(\mathbf{R}_{x}).$
Note that $\widetilde{\mathbf{R}}_{\tilde{s}}$ and $f(\mathbf{R}_{x})$ are both the linear functions w.r.t. variable $\mathbf{R}_{x}$ with a given point $\mathbf{R}_{0}$. By substituting (54) and (55) into $\mathcal{P}_{2}$, the problem can be relaxed into
$\mathcal{P}_{3}\left\{\begin{aligned} \mathop{\min}\limits_{\mathbf{R}_{x},\mathbf{D}\in\mathbb{S}_{N_{t}}^{+}}&D_{s}(\mathbf{R}_{x})+\text{Tr}\left(\mathbf{D}\right)\\ \text{subject to}&\kern 5.0ptC(\mathbf{R}_{x})-M_{s}\left(f(\mathbf{R}_{x})-\log\det(\mathbf{D})\right)\geq 0,\\ &\kern 5.0pt\widetilde{\mathbf{R}}_{\tilde{s}}-\mathbf{D}\succeq 0,\kern 2.0pt\mathbf{R}_{x}\succeq 0,\kern 2.0pt\text{Tr}\left(\mathbf{R}_{x}\right)\leq P_{T}.\end{aligned}\right.$ (56)
For a given point $\mathbf{R}_{0}$, $\mathcal{P}_{3}$ is convex since both the objective and constraints are either convex or linear. Therefore, it can be efficiently solved by using the off-the-shelf CVX toolbox [35]. For an initial covariance matrix, e.g., $\mathbf{R}_{0}=P_{T}/N_{t}\mathbf{I}$, the non-convex problem $\mathcal{P}_{2}$ can be addressed by iteratively computing the optimal solution $\mathbf{R}^{\star}_{x}$ of $\mathcal{P}_{3}$ and updating $\mathbf{R}_{0}=\mathbf{R}^{\star}_{x}$ until convergence. The procedure is outlined in Algorithm 4.
VI Simulation Results
The simulation results are divided into two sections. The first section evaluates the optimal distribution for the ISAC channel input in scalar cases, demonstrating the effectiveness of the proposed MSST and the two-step BA-based optimal search algorithm. Moreover, the unique distortion-capacity (D-C) curves of the CAS system are provided to illustrate the performance tradeoff between S&C processes. The second section presents the results of the optimal waveform design for the Gaussian MIMO channel, showing its superiority over the S&C-optimal schemes.
VI-A Optimal Distribution for ISAC Channel Input
In this subsection, we aim to reveal the unique performance tradeoff between S&C processes. The simulation parameters are set as follows. The prior distribution of the target’s sate is assumed to be $P_{\mathsf{S}}(s)\sim\mathcal{CN}(0,1)$, i.e, $\nu_{s}^{2}=1$. We define the normalized signal-to-noise ratios (SNRs) for the S&C channels as $10\log(1/\sigma^{2})$ dB and set the power budget to be $B=5$. Unless otherwise specified, the S&C channel SNRs are set to be $\text{SNR}_{s}=\text{SNR}_{c}=0$ dB in this subsection. To improve the computational efficiency, the weighted factor $\mu$ takes values from the sets $[0,1]$, $[1,5]$, and $[5,30]$ with non-uniform spacings of $0.1$, $0.5$, and $5$, respectively.
Figure 3: The changes of the S&C distortions with various factor $\mu$.
Figure 4: The CAS distortion versus the communication channel capacity.
Figure 5: The channel input distribution of S&C-optimal and CAS-optimal.
VI-A1 The D-C curves and the optimal distribution of ISAC channel input
Fig. 3 presents the D-C curves for both S&C processes. As expected, sensing distortion $D_{s}$ increases as the values of $\mu$ decrease, implying that the estimation performance at the BS becomes less significant. Conversely, the communication channel capacity increases accordingly, leading to a reduction in communication distortion $D_{c}$. Consequently, these opposing trends in S&C distortions result in an uncertain CAS distortion. In Fig. 4, we show the resultant D-C curve for the CAS systems with different sensing SNRs. A CAS-optimal point that minimizes the CAS distortion is evident, highlighting a unique phenomenon in CAS systems compared to the existing studies [14, 15].
Fig. 5 depicts the specific channel input distributions corresponding to three special points in Fig.4, namely, the S&C-optimal and CAS-optimal points. The channel capacity is maximized without the sensing constraint when $\mu=0$. It is well-known that Gaussian channel input achieves the maximum MI for the Gaussian channels. By contrast, at the sensing-optimal point, as $\mu$ becomes sufficiently large, the channel input exhibits a 2-ary pulse amplitude modulation. These results are consistent with the findings in [12]. However, the unique CAS-optimal distribution, presenting a compromise between the S&C-optimal distributions, is attained at $\mu=2.5$.
Figure 6: The distortion-capacity curves with various state variances.
Figure 7: The CAS-optimal distribution for various state variances.
VI-A2 The impact of the target’s prior variance
Subsequently, we investigate the impact of the target’s state variance by selecting the values of $\nu_{s}^{2}\in\{0.1,0.2,0.5,5\}$. Due to the significant differences in the order of magnitude of the values for MI and sensing distortions in (32), we adjust the interval of $\mu$ for each state variance rather than using a fixed interval. The essential purpose of the CAS systems is to reduce the uncertainty of the target’s prior information, specifically the variance of the state’s prior distribution.
In Fig. 6, each point represents a lower CAS distortion compared to the prior variance, even at the S&C-optimal points. This reduction is attributed to the resource multiplexing gain achieved through the dual-functional signaling strategy. Additionally, we observe that the CAS-optimal point gradually shifts from sensing-optimal to communication-optimal as variance increases. A small variance indicates relatively accurate prior information about the target’s state, necessitating an ISAC channel input with strong sensing capabilities for better accuracy performance. Conversely, as variance increases, greater emphasis on communication performance is required to convey more information.
We provide the detailed distributions of the CAS-optimal channel input for different state variances. For a small variance (e.g., $\nu_{s}^{2}=0.1$), the sensing distortion becomes small accordingly. It is at least less than the prior variance; otherwise, the user can ‘guess’ the state information in terms of the prior distribution. Therefore, a large weighted factor is required, e.g., $\mu=690$, to balance the values of MI and sensing distortion in (32). In this case, the CAS-optimal distribution aligns with the sensing-optimal distribution, specifically a standard 2-ary pulse amplitude modulation. As anticipated, the optimal $\mu$ values decrease with increasing state variance, and the CAS-optimal distribution changes accordingly.
VI-B Optimal Waveform Design for Gaussian Signal
In this subsection, we demonstrate the effectiveness of the proposed ISAC waveform design scheme for Gaussian signaling. The general system setups are as follows. The communication channel is modeled by Rayleigh fading, where each entry of $\mathbf{H}_{c}$ obeys the standard complex Gaussian distribution. All results are obtained by the average of $100$ Monte Carlo trials and the average distortion in each Monte Carlo trial is calculated by $D/(M_{s}N_{t})$.
Figure 8: The average distortion versus sensing channel SNRs for various methods, $a_{1}=0.4,a_{2}=0.1$, $N_{t}=M_{c}=M_{s}=2$.
VI-B1 Waveform design for a 2-D special case
We begin with a two-dimensional (2-D) special case where the state covariance matrix, $\bm{\Sigma_{s}}=\text{diag}\{a_{1},a_{2}\}$, is assumed to be diagonal. In this scenario, the global optimal solution of problem $\mathcal{P}_{2}$ can be attained through a 2-D exhaustive search, as outlined in [11]. Thus, this optimal solution  [^4] The optimal eigenspace of $\mathbf{R}_{x}$ must align with the communication eigenspace of $\mathbf{H}_{c}^{H}\mathbf{H}_{c}$ as the sensing eigenspace is an arbitrary unitary matrix. Thus, the original problem can be reduced to finding two optimal eigenvalues over 2-D grids, with the sum of each 2-D point equal to $P_{T}$. can serve as a benchmark for comparing the superiority of our method. Accordingly, the numbers of the Tx and S&C Rx antennas are set by $N_{t}=M_{c}=M_{s}=2$.
Fig. 8 illustrates the impact of sensing channel quality on the achievable average distortion. The sensing SNRs vary from 00 dB to $20$ dB with a spcing $2$ dB. Besides the global optimal benchmark, the S&C optimal waveform schemes refer to the solutions that minimize the MMSE in (49) and maximize the channel capacity in (52), respectively. At a glance, we observe that the achievable average distortion decreases with increasing sensing channel SNRs. Furthermore, the sensing-optimal scheme is preferable for acquiring accurate target information when the communication channel is sufficiently good at $\text{SNR}_{c}=5$ dB. However, the sensing-optimal scheme may cause significant performance degradation as the communication SNR decreases. Particularly, the sensing-optimal curve even tends to be flat at $\text{SNR}_{c}=-5$ dB, implying that a poor communication channel severely limits performance improvement. This clearly highlights the unique tadeoff between S&C processes in CAS systems. Finally, it is worth noting that the proposed algorithm achieves satisfactory performance compared to the global optimal solution.
Figure 9: The impact of the state variances on the CAS performance, $\text{SNR}_{s}=5$ dB, $N_{t}=M_{c}=M_{s}=2$.
In Fig. 9, we investigate the impact of state variances on the achievable average distortion in the MIMO systems. The state variances are scaled by a factor ranging from $2$ to $10$. Two interesting observations emerge. 1) For small state variances, the global optimal scheme coincides with the sensing-optimal scheme at $\text{SNR}_{c}=10$ dB. However, both the S&C-optimal schemes experience performance degradation as state variances increase. 2) When the communication channel quality is insufficient at $\text{SNR}_{c}=-5$ dB, the global optimal scheme tends towards the communication-optimal scheme to enhance communication capabilities. As state variances increase, the performance degradation of the sensing-optimal scheme becomes more significant. Nevertheless, our proposed algorithm can effectively balance the S&C processes, resulting in improved average distortion.
Figure 10: The average distortion versus sensing channel SNRs for various methods, $N_{t}=10,M_{s}=2,M_{c}=5$.
VI-B2 Waveform design for general MIMO cases
Subsequently, we extend the results to high-dimensional matrix scenarios, where the number of Tx antennas is $N_{t}=10$, and the numbers of S&C Rx antennas are $M_{s}=2$ and $M_{c}=5$, respectively. The state covariance matrix $\bm{\Sigma}_{s}$ is randomly generated as an $N_{t}\times N_{t}$ Hermitian matrix. It is worth noting that obtaining the global optimal solution is challenging due to the highly coupled S&C processes. For comparison purposes, we provide the results of the Heuristic algorithm proposed in [11], whose formulation is given by
$\mathop{\max}\limits_{\mathbf{R}_{x}}$ $\kern 5.0pt\beta I(\mathbf{Y};\mathbf{X}|\mathbf{H}_{c})+(1-\beta)I(\mathbf{Z};\mathbf{H}_{s}|\mathbf{X})$ (57)
subject to $\kern 5.0pt\text{Tr}\left(\mathbf{R}_{x}\right)\leq P_{T},$
where $I(\mathbf{Z};\mathbf{H}_{s}|\mathbf{X})=\log\det\left(\frac{1}{\sigma^{2}_{s}}\bm{\Sigma}_{s}\mathbf{R}_{x}+\mathbf{I}_{N_{t}}\right)$ is defined by sensing MI and $I(\mathbf{Y};\mathbf{X}|\mathbf{H}_{c})$ is defined in (52). The weighted factor $\beta$ takes values over the interval $[0,1]$ with $L$ grid points. In our simulations, $L$ is set to be $11$.
Fig. 10 illustrates the average distortions versus sensing channel SNRs for various waveform design schemes. It can be observed that the proposed algorithm outperforms its counterparts. Similar trends to those shown in Fig. 8 for the 2-D case are also noticeable. Besides, we can observe that the proposed ISAC waveform design achieves significantly greater performance gains in high-dimensional scenarios.
Figure 11: The average distortion versus the number of sensing Rx antennas, $N_{t}=10,M_{c}=5$, $\text{SNR}_{s}=0$ dB, $\text{SNR}_{c}=10$ dB.
Figure 12: The average distortion versus the number of communication Rx antennas, $N_{t}=10,M_{s}=2$, $\text{SNR}_{c}=0$ dB.
Fig. 11 and Fig. 12 demonstrate the influence of S&C Rx antennas on average distortion across various waveform design schemes. Fig. 11 exhibits a linear increase in average distortion with the addition of sensing Rx antennas. The linearity feature consistent with the Kronecker product form in assumption 2. Moreover, regarding the TRM estimation, an increase in sensing Rx antennas necessitates more channel information to be estimated, leading to a larger average distortion.
In contrast, Fig. 12 shows a declining trend in average distortion as the number of communication Rx antennas increases. This performance improvement is attributed to enhanced communication channel capacity facilitated by multiple Rx antenna gain, thereby improving overall system performance. Additionally, under the communication-optimal scheme, there is a gradual performance degradation as channel capacity increases. This trend arises because improving sensing performance becomes critical in scenarios with excessive communication channel capacity. Once again, the proposed ISAC waveform design method is superior to the other counterparts in all scenarios.
VII Conclusion
In this paper, we delve into the dual-functional signaling strategy for communication-assisted sensing (CAS) systems across three primary aspects. Firstly, we develop a modified source-channel separation theorem (MSST) tailored for CAS systems under the condition of separable distortion metric. We elucidate the operational meaning of the proposed MSST from the perspective of coding theory. Secondly, we develop an input distribution optimization scheme for the CAS system under scalar channels, which minimizes the CAS distortion while adhering to the MSST and resource constraints. We determine the optimal distribution for CAS systems, balancing between the conventional communication-optimal (Gaussian distribution) and sensing-optimal (2-ary pulse amplitude modulation) schemes. Thirdly, we propose an ISAC waveform design method for Gaussian signaling in MIMO CAS systems, where a successive convex approximation (SCA) algorithm is conceived to solve for the optimal covariance matrix. Simulation results show the unique performance tradeoff between the sensing and communication processes.